# Cutwidth Bounds via Vertex Partitions

### v1

@@ Abstract
We study the cutwidth measure on graphs and ways to bound the cutwidth of a graph by partitioning its vertices. We consider bounds expressed as a function of two quantities: on the one hand, the maximal cutwidth $x$ of the subgraphs induced by the classes of the partition, and on the other hand, the cutwidth $y$ of the quotient multigraph obtained by merging each class to a single vertex. We consider in particular decomposition of directed graphs into strongly connected components (SCCs): in this case, $x$ is the maximal cutwidth of an SCC, and $y$ is the cutwidth of the directed acyclic condensation multigraph.
@@ 1 Introduction
In this paper, we study how to show upper bounds on the treewidth of graphs via the approach of partitioning its vertices into classes, and studying on the one hand the cutwidth of each class, and on the other hand the cutwidth of the quotient obtained by merging each class to a single vertex. As the quotient may in general feature parallel edges, and cutwidth accounts for these parallel edges, then we define the quotient as a multigraph, and work with multigraphs throughout the paper. Let us assume that we can show an upper bound of $x$ on the cutwidth of the quotient multigraph, and that we can show an upper bound of $y$ on the maximal cutwidth of a class of partition. We would then want to achieve a bound on the cutwidth of the original graph, as a function of $x$ and $y$. Intuitively, we want to do so by ordering the classes according to the order that achieves cutwidth $x$ on the quotient multigraph, and then ordering each class according to the order that achieves cutwidth at most $y$ within each class.
This technique of bounding cutwidth via vertex partitions has been studied earlier. barth1995bandwidth have studied how to bound the bandwidth or cutwidth of a graph as a function of the same measure on the quotient graph and on the classes of the partition; but their bounds depend on the degree of the graph. More recently, FeldmannM23 have applied the method in the specific case where the input graph is directed, and the classes of the partition are the strongly connected components (SCCs): the quotient multigraph is then called the condensation multigraph. Their work (FeldmannM23, Lemma 2.2) claims an upper bound of $x+y$ on the cutwidth of the graph $G$, again as a function of the (undirected) cutwidth $x$ of the condensation multigraph of $G$ and of the maximal (undirected) cutwidth $y$ of an SCC of $G$. This standalone lemma of FeldmannM23 is of independent interest, and we recently applied it to our study of the edge-minimum walk of modular length problem in amarilli2025edge.
@@
Our result achieves a cutwidth bound $1.5x+y$ and not $x+y$, but as we show, this is unavoidable (and the bound of $x+y$ claimed in (FeldmannM23, Lemma 2.2) is in fact not correct). Specifically, we show in Proposition 5.1 that, for arbitrarily large $x$ and $y$, there are graphs whose cutwidth is $1.5x+y$: this implies that the bound of $1.5x+y$ cannot be improved in general.

### v2

@@ Abstract
We study the cutwidth measure on graphs and ways to bound the cutwidth of a graph by partitioning its vertices. We consider bounds expressed as a function of two quantities: on the one hand, the maximal cutwidth $y$ of the subgraphs induced by the classes of the partition, and on the other hand, the cutwidth $x$ of the quotient multigraph obtained by merging each class to a single vertex. We consider in particular the decomposition of directed graphs into strongly connected components (SCCs): in this case, $y$ is the maximal cutwidth of an SCC, and $x$ is the cutwidth of the directed acyclic condensation multigraph.
@@ 1 Introduction
In this paper, we study how to show upper bounds on the treewidth of graphs via the approach of partitioning its vertices into classes, and studying on the one hand the cutwidth of each class, and on the other hand the cutwidth of the quotient obtained by merging each class to a single vertex. As the quotient may in general feature parallel edges, and cutwidth accounts for these parallel edges, then we define the quotient as a multigraph, and work with multigraphs throughout the paper. Let us assume that we can show an upper bound of $x$ on the cutwidth of the quotient multigraph, and that we can show an upper bound of $y$ on the maximal cutwidth of the classes of the partition. We would then want to show an upper bound on the cutwidth of the original graph, as a function of $x$ and $y$. Intuitively, we want to do so by ordering the classes according to the order that achieves cutwidth $x$ on the quotient multigraph, and then ordering each class according to the order that achieves cutwidth at most $y$ within each class.
This technique of bounding cutwidth via vertex partitions has been studied earlier. barth1995bandwidth have studied how to bound the bandwidth or cutwidth of a graph as a function of the same measure on the quotient graph and on the classes of the partition; however, their bounds depend on the degree of the graph. More recently, FeldmannM23 have applied the method in the specific case where the input graph is directed, and the classes of the partition are the strongly connected components (SCCs): the quotient multigraph is then called the condensation multigraph. Their work (FeldmannM23, Lemma 2.2) claims an upper bound of $x+y$ on the cutwidth of the graph $G$, again as a function of the (undirected) cutwidth $x$ of the condensation multigraph of $G$ and of the maximal (undirected) cutwidth $y$ of an SCC of $G$. This standalone lemma of FeldmannM23 is of independent interest, and we recently applied it to our study of the edge-minimum walk of modular length problem in amarilli2025edge.
@@
Our result achieves a cutwidth bound of $1.5x+y$ and not $x+y$, but as we show, this is unavoidable (and the bound of $x+y$ claimed in (FeldmannM23, Lemma 2.2) is in fact not correct). Specifically, we show in Proposition 5.1 that, for arbitrarily large $x$ and $y$, there are graphs whose cutwidth is $1.5x+y$: this implies that the factor $1.5$ cannot be improved in general.